Brett has 2 gallons of 35% antifreeze. How many gallons of 100% antifreeze should he add to the solution to produce a solution of 50% antifreeze?
Found 2 solutions by josgarithmetic, richwmiller:
Answer by josgarithmetic(39617)
Answer by richwmiller(17219)

1*x+0.35*2=0.5(2+x)
1*x+0.7=1+0.5x
1x-0.5x=1-0.7
0.5x=0.3
x=0.6 gal of 100% is added to 2 gal of 35% to get 2.6 gal at 50%
check
1*0.6+0.35*2=0.5(2+0.6)
0.6+0.7=0.5(2.6)
1.3=1.3
ok